tom's bike shop rents bikes for $0.30 a mile and an initial $10.00 fee. joe's bike shop rents bikes for $0.50 a mile and an initial fee of $6.00. at what mileage would the two shops cost the same? what would that cost be?
Found 2 solutions by mananth, checkley77:
Answer by mananth(16949)

tom's bike shop rents bikes for $0.30 a mile and an initial $10.00 fee. joe's bike shop rents bikes for $0.50 a mile and an initial fee of $6.00. at what mileage would the two shops cost the same? what would that cost be?

let the number of miles be x
TOM
0.3x++10
JOE
0.5x+6
..
0.3x+10=0.5x+6
-0.3x
0.3x-0.3x+10=0.5x-0.3x+6
10=0.2x+6
-6
10-6=0.2x
4=0.2x
x=4/0.2
= 20 miles.
At 20 miles both will cost the same

Answer by checkley77(12844)

10+.30x=6+.50x
.30x-.50x=6-10
-.20x=-4
x=-4/-.20
x=20 miles makes the 2 deals equal.
10+.30*20=10+6=$16:00 is the cost of each offer for 30 miles.
6+.50*20=6+10=$16:00 ditto.